OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2= 8x-35

OpenStudy (anonymous):

x^2-8x+35=0 try to solve it and i will check it after u

OpenStudy (anonymous):

@GreenPride

OpenStudy (anonymous):

quadratic formula \[\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]\[a = 1\]\[b = -8\]\[c = 35\]

OpenStudy (whpalmer4):

To use the quadratic formula, get your equation in the proper form: \[ax^2 + bx + c = 0,\,a\ne 0\]@Abdulhameed has already done this for you. Then plug the numbers into the formula \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]You'll get two solutions unless \(b^2-4ac = 0\) (a perfect square)

OpenStudy (anonymous):

When I plug it in I get a different number than 19 :/ And the answer is \[4+i \sqrt{19} and 4 - i \sqrt{19}\]

OpenStudy (anonymous):

\[\frac{ -8 \pm \sqrt{64-140} }{ 2 }\]\[\frac{ -8 \pm \sqrt{-76} }{ 2 }\]\[-4 \pm \frac{ i \sqrt{76} }{ 2 }\]

OpenStudy (anonymous):

Thats it ?

OpenStudy (anonymous):

yeah, can't be simplify it anymore

OpenStudy (whpalmer4):

@helpme1.2 that's not true, you can simplify it to \[4\pm i\sqrt{19}\] \[-4\pm \frac{i\sqrt{76}}2 = -4\pm\frac{i\sqrt{2*2*19}}{2} = -4\pm\frac{i*2*\sqrt{19}}2=-4\pm i\sqrt{19}\]

OpenStudy (whpalmer4):

Sorry, that should be \[4\pm i\sqrt{19}\]I copied your result, which has an incorrect sign because you didn't change the sign of \(b\) in the fraction.

OpenStudy (whpalmer4):

\[a = 1,\, b = -8,\, c = 35\]\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-(-8) \pm\sqrt{(-8)^2-4(1)(35)}}{2(1)} \]\[=\frac{8\pm\sqrt{64-140}}{2} = 4\pm\frac{\sqrt{-76}}2 = 4\pm i\sqrt{19}\]